Jefferson Lab HALL A Logbook for September 06 - All EXCEPT Automatic Entries

  • Main INDEX

    Saturday

  • 186024 09/30/06 23:59 xiaohui Target Screen
  • 186021 09/30/06 23:06 Fatiha HallA tool screen for the new kinematics (Q2=0.3189).
  • 186020 09/30/06 23:04 Yi Qiang Spot++ check
  • 186018 09/30/06 22:28 xiaohui Target: 15cm LH2
  • 186017 09/30/06 22:18 atrig Trigger-Download
  • 186016 09/30/06 22:17 atrig Trigger-Download
  • 186015 09/30/06 22:01 atrig Trigger-Download
  • 186014 09/30/06 21:59 atrig Trigger-Download
  • 186013 09/30/06 21:56 Fatiha Status
  • 186012 09/30/06 21:53 atrig Trigger-Download
  • 186009 09/30/06 20:39 xiaohui Target: 15cm Dummy
  • 186006 09/30/06 20:15 xiaohui Target: Carbon
  • 186002 09/30/06 19:29 Fatiha Bad quality beam since 25 min
  • 185997 09/30/06 18:11 Fatiha Fast Feed-Back off
  • 185995 09/30/06 18:02 Fatiha Beam Feed back, positionm etc...
  • 185994 09/30/06 17:55 Yi Qiang BPM Y Profile through Run 5386
  • 185992 09/30/06 17:39 Fatiha Fast feed back is ON now and spot shape much better! :)
  • 185991 09/30/06 17:38 Yi Qiang FB ON solved the BPM puzzle!!
  • 185988 09/30/06 17:31 Fatiha Fast Feed Back is off since???
  • 185985 09/30/06 16:58 Yi Qiang Beam postion Y profile of Run 5385
  • 185984 09/30/06 16:57 Fatiha Added the missing angle information in the Hall A general tool
  • 185981 09/30/06 16:39 Fatiha Magnet and current snap shots
  • 185979 09/30/06 16:02 Kalyan Allad Shift/Run Summary
  • 185977 09/30/06 15:55 kalyan no beam for last 25mins
  • 185976 09/30/06 15:32 kalyan tech problem with RF seperator
  • 185968 09/30/06 13:27 B. Craver spot++
  • 185967 09/30/06 13:27 B. Craver spot is stable
  • 185966 09/30/06 13:26 kalyan beam is back
  • 185964 09/30/06 12:42 kalyan Q1 tripped again on both left and right arm
  • 185963 09/30/06 12:27 S. Strauch HRS magnet cryo-level problems
  • 185962 09/30/06 12:26 B. Craver beam spot plots
  • 185961 09/30/06 12:25 kalyan still no beam!
  • 185960 09/30/06 11:13 kalyan ready to take the beam
  • 185959 09/30/06 10:39 kalyan Hall in controlled access
  • 185958 09/30/06 10:35 B. Craver target stripcharts
  • 185957 09/30/06 10:34 kalyan asked for controlled access
  • 185956 09/30/06 10:17 kalyan strip chart for He levels
  • 185955 09/30/06 10:09 kalyan beam position chart
  • 185954 09/30/06 10:08 kalyan Q1 tripped
  • 185952 09/30/06 09:54 jglister spot++ still abnormal
  • 185949 09/30/06 08:57 kalyan beam is back
  • 185947 09/30/06 08:25 kalyan no beam
  • 185944 09/30/06 08:00 G.M. Urciuol 9/30/2006 owl shift report
  • 185941 09/30/06 06:56 G.M. Urciuol HRS Left Arm Momentum Change
  • 185932 09/30/06 03:59 Elena Kh. Target screens.
  • 185927 09/30/06 01:20 G.M. Urciuol Q1 of left and rigth arm now o.k.
  • Friday

  • 185925 09/29/06 23:54 A. Puckett magnet cryogen levels recovering
  • 185924 09/29/06 23:33 A. Puckett Cryogen levels crashed on several Left and Right HRS magnets
  • 185912 09/29/06 18:14 A. Puckett Helium levels in Left Dipole uncomfortably low
  • 185904 09/29/06 16:04 R. Michaels scan of T3, and coinc. time
  • 185903 09/29/06 15:57 B. Craver Day Shift Summary
  • 185901 09/29/06 15:33 atrig Trigger-Download
  • 185884 09/29/06 15:09 atrig Trigger-Download
  • 185879 09/29/06 14:09 X. Jiang rate at ep kin-3.
  • 185877 09/29/06 13:31 LeRose and more target charts
  • 185876 09/29/06 13:28 LeRose More Target charts
  • 185875 09/29/06 13:27 LeRose target charts
  • 185874 09/29/06 13:22 B. Craver Increase current
  • 185873 09/29/06 13:20 R. Michaels/ coinc. time
  • 185870 09/29/06 13:12 B. Craver Coincidence trigger
  • 185868 09/29/06 13:04 R. Michaels trigger access activity
  • 185867 09/29/06 12:58 B. Craver beam permit
  • 185866 09/29/06 12:54 atrig Trigger-Download
  • 185863 09/29/06 12:13 atrig Trigger-Download
  • 185862 09/29/06 12:12 atrig Trigger-Download
  • 185861 09/29/06 12:11 atrig Trigger-Download
  • 185860 09/29/06 12:03 atrig Trigger-Download
  • 185859 09/29/06 11:51 atrig Trigger-Download
  • 185858 09/29/06 11:50 atrig Trigger-Download
  • 185857 09/29/06 11:46 atrig Trigger-Download
  • 185856 09/29/06 11:42 atrig Trigger-Download
  • 185854 09/29/06 11:11 atrig Trigger-Download
  • 185853 09/29/06 10:59 atrig Trigger-Download
  • 185852 09/29/06 10:39 B. Craver Controlled Access
  • 185851 09/29/06 10:38 R. Michaels access to figure out coinc.
  • 185845 09/29/06 09:51 B. Craver beam position unstable?
  • 185844 09/29/06 09:49 R. Michaels run 5341-43 suspect quality
  • 185841 09/29/06 09:34 R. Michaels edtmsetup wrong ?
  • 185835 09/29/06 08:48 atrig Trigger-Download
  • 185833 09/29/06 08:46 B. Craver reduce beam current
  • 185832 09/29/06 08:45 B. Craver Beam calorimeter calibration
  • 185827 09/29/06 08:24 gron Split file in CODA
  • 185826 09/29/06 08:17 R. Michaels trigsetup GUI fix
  • 185822 09/29/06 08:00 G.M. Urciuol Target Halog entry for 9/29/2006 owl shift
  • 185821 09/29/06 07:57 paschke Owl Shift Run Summary (re-formatted)
  • 185820 09/29/06 07:56 paschke Owl Shift Summary
  • 185819 09/29/06 07:54 paschke Owl Shift Run Summary
  • 185818 09/29/06 07:39 paschke Trigger delay twiddled, restored.
  • 185816 09/29/06 07:33 atrig Trigger-Download
  • 185815 09/29/06 07:25 atrig Trigger-Download
  • 185814 09/29/06 07:05 atrig Trigger-Download
  • 185810 09/29/06 06:22 S. Frullani VDC left restored
  • 185807 09/29/06 05:56 S. Frullani hot wire in VDC left
  • 185806 09/29/06 05:42 paschke Coincidence trigger delay range understood, but still not working
  • 185804 09/29/06 05:27 atrig Trigger-Download
  • 185803 09/29/06 05:26 atrig Trigger-Download
  • 185793 09/29/06 02:34 paschke coincidence trigger not working
  • 185791 09/29/06 02:27 G.M. Urciuol Target moved to H2 15 cm
  • 185789 09/29/06 02:15 paschke Coincidence trigger problem: delay out of range
  • 185787 09/29/06 02:11 G.M. Urciuol Target moved to 15 cm dummy
  • 185782 09/29/06 01:44 atrig Trigger-Download
  • 185780 09/29/06 01:18 G.M. Urciuol Target moved to 15 cm dummy and to carbon
  • 185772 09/29/06 00:02 A. Puckett Swing Shift Summary
  • Thursday

  • 185766 09/28/06 23:13 mpaolone cryotarg screen
  • 185765 09/28/06 23:13 A. Puckett ioc reboot howto, continued
  • 185764 09/28/06 23:05 A. Puckett Quick HOWTO of IOC reboot
  • 185763 09/28/06 22:58 A. Puckett lessons learned from magnet crash
  • 185754 09/28/06 20:32 A. Puckett Back up and running
  • 185752 09/28/06 20:13 A. Puckett left magnet IOC reboot
  • 185751 09/28/06 19:54 A. Puckett rebooted left magnet IOCs
  • 185750 09/28/06 19:19 A. Puckett Left Arm NMR not working
  • 185749 09/28/06 19:09 A. Puckett Lost Left Arm Dipole, Reset.
  • 185745 09/28/06 17:52 eip run 5314-beam calibartion
  • 185738 09/28/06 16:26 R. Michaels helicity ok for split files
  • 185734 09/28/06 16:05 Ole Hansen Shift Summary
  • 185733 09/28/06 16:01 X. Jiang FPP door. 3/4 carbon is IN, plus 3" door.
  • 185732 09/28/06 16:00 X. Jiang FPP threshold at run 5310.
  • 185731 09/28/06 15:59 X. Jiang FPP HV low-V at run 5310
  • 185730 09/28/06 15:57 X. Jiang magnet screen at run 5310.
  • 185715 09/28/06 14:52 R. Michaels Helicity check du jour
  • 185709 09/28/06 14:25 Ole Hansen delta = 0
  • 185699 09/28/06 13:27 Ole Hansen Beam current history
  • 185698 09/28/06 13:26 Ole Hansen FPP panels
  • 185697 09/28/06 13:25 Ole Hansen Hall A Tools
  • 185694 09/28/06 12:41 jglister spot++
  • 185692 09/28/06 12:17 Elena Kh. Target screens.
  • 185678 09/28/06 11:08 Ole Hansen Helicity problems?
  • 185674 09/28/06 09:14 X. Jiang run 5286 taken during the calibration of Siver calorimeter.
  • 185673 09/28/06 09:06 Ole Hansen HRSR to 61.43 degrees
  • 185672 09/28/06 09:06 Ole Hansen NMR fixed
  • 185671 09/28/06 08:48 Ole Hansen HRSR dipole NMR frozen
  • 185670 09/28/06 08:44 Ole Hansen HRSR momentum change
  • 185669 09/28/06 08:43 Ole Hansen Access
  • 185665 09/28/06 08:07 G.M. Urciuol 9/28/2006 owl shift TO Halog entry
  • 185663 09/28/06 07:58 paschke Owl Shift Run Summary
  • 185662 09/28/06 07:57 paschke Owl Shift Summary (reformatted)
  • 185661 09/28/06 07:56 paschke Owl Shift Summary
  • 185652 09/28/06 05:10 paschke spot weird again
  • 185642 09/28/06 02:44 G.M. Urciuol Target moved to H2 15 cm
  • 185639 09/28/06 02:27 G.M. Urciuol Target moved to Carbon
  • 185636 09/28/06 02:11 G.M. Urciuol Tsrget Moved to 15 cm dummy
  • Wednesday

  • 185625 09/27/06 23:50 A. Puckett Swing Shift Summary
  • 185616 09/27/06 20:43 mpaolone Cryotarg Screen
  • 185612 09/27/06 20:17 A. Puckett Beamline high voltages missing... problem?
  • 185601 09/27/06 17:13 atrig Trigger-Download
  • 185599 09/27/06 16:48 R. Michaels Switched to 1-daq mode
  • 185595 09/27/06 16:26 atrig Trigger-Download
  • 185593 09/27/06 15:54 jpchen controlled access status
  • 185592 09/27/06 15:37 gilman FPP carbon door
  • 185591 09/27/06 15:08 jpchen day shift summary
  • 185590 09/27/06 15:08 jpchen planned controlled access activities.
  • 185589 09/27/06 15:02 jpchen ED data taking completed. requested controlled access for configuration change
  • 185584 09/27/06 14:28 Elena Kh. 4 cm Dummy target
  • 185571 09/27/06 13:40 Elena Kh. LH2, 4 cm target
  • 185558 09/27/06 12:24 Elena Kh. LD2, 4 cm target
  • 185555 09/27/06 12:12 Ole Hansen Beam position drift
  • 185552 09/27/06 12:00 Elena Kh. Target screens
  • 185551 09/27/06 11:58 Elena Kh. Target screens
  • 185550 09/27/06 11:57 Elena Kh. Ta target
  • 185547 09/27/06 11:49 Ole Hansen spot++ results
  • 185546 09/27/06 11:48 Ole Hansen Beam spot size and position
  • 185534 09/27/06 09:58 jpchen Controlled access finished. Requested beam
  • 185533 09/27/06 09:46 jpchen right nmr communication problem fixed. Spectrometer all set.
  • 185532 09/27/06 09:42 jpchen 1.5" carbon door closed
  • 185531 09/27/06 09:11 jpchen Configuration change to kin near completion, rigght dipole NMR problem
  • 185530 09/27/06 08:31 paschke Owl Shift Summary
  • 185528 09/27/06 08:02 paschke Owl Shift Run Summary
  • 185527 09/27/06 07:59 paschke White Spectrum Study
  • 185526 09/27/06 07:42 G.M. Urciuol 9/27/2006 owl shift
  • 185519 09/27/06 07:13 paschke Tools screen for white #6
  • 185514 09/27/06 06:44 paschke Tool screen for white point #5
  • 185505 09/27/06 05:57 paschke Hall A Tools for white point #3
  • 185488 09/27/06 04:56 paschke Hall A Tools for white setting #2
  • 185477 09/27/06 03:27 G.M. Urciuol Target moved
  • 185474 09/27/06 03:09 paschke playing with position lock fixed spot
  • 185467 09/27/06 02:56 G.M. Urciuol Tasrget moved
  • 185456 09/27/06 02:27 G.M. Urciuol Alarm and a limit changed
  • 185451 09/27/06 01:56 paschke spot++ still abnormal
  • 185436 09/27/06 01:10 G.M. Urciuol Target moved to H2 4cm
  • 185433 09/27/06 00:38 paschke Hall A Tools screenshot
  • 185432 09/27/06 00:27 A. Camsonne HV for PMT
  • 185431 09/27/06 00:23 A. Camsonne Compton FADC test with Hamatsu PMT
  • 185430 09/27/06 00:04 A. Puckett Swing Shift Summary
  • Tuesday

  • 185424 09/26/06 23:06 mpaolone Cryotarg Screen
  • 185423 09/26/06 22:53 mpaolone Target Change to LD2
  • 185422 09/26/06 22:41 A. Puckett Setting left HRS dipole
  • 185415 09/26/06 21:09 mpaolone Low H2 tank pressure alarm
  • 185412 09/26/06 21:01 mpaolone Target change to 4cm dummy
  • 185386 09/26/06 16:02 Ole Hansen Shift Summary
  • 185383 09/26/06 15:46 Elena Kh. 4 cm Dummy target
  • 185368 09/26/06 11:20 jglister spot++ plot
  • 185367 09/26/06 11:15 jglister spot++
  • 185366 09/26/06 11:10 Ole Hansen Beam current history
  • 185365 09/26/06 11:10 Ole Hansen Hall A Tools
  • 185364 09/26/06 11:09 M Jones Raised Front FPP HV from 1775 to 1800
  • 185363 09/26/06 11:06 Elena Kh. Target screens
  • 185356 09/26/06 09:28 Elena Kh. LD2, 4 cm target
  • 185352 09/26/06 08:02 paschke Owl Shift Summary
  • 185351 09/26/06 08:01 paschke Owl Shift Run Summary
  • 185348 09/26/06 07:13 kelleher target screens captured!
  • 185347 09/26/06 07:11 kelleher target motion: C
  • 185342 09/26/06 05:25 paschke kinematics 12c summary: various beam currents used
  • 185341 09/26/06 05:10 kelleher target motion: Ta
  • 185336 09/26/06 04:33 kelleher target motion: 4cm Dummy
  • 185333 09/26/06 04:22 paschke spot++ seems abnormal
  • 185318 09/26/06 03:25 kelleher target motion: hydrogen
  • 185309 09/26/06 02:13 kelleher target motion: Al, the target position
  • Monday

  • 185295 09/25/06 23:58 Xin Qian Target Checklist
  • 185288 09/25/06 22:50 A. Puckett Swing Shift Summary
  • 185287 09/25/06 22:33 A. Puckett controlled access to move ladder before HRSL angle change
  • 185286 09/25/06 22:09 A. Puckett HRSL angle change
  • 185281 09/25/06 20:45 A. Puckett moved target to 4 cm Al dummy cell for kinematics 11c
  • 185256 09/25/06 19:13 A. Puckett changed HRSL momentum setting.
  • 185255 09/25/06 19:10 A. Puckett moved target to LH2
  • 185252 09/25/06 18:59 X. Jiang magnet screen at kinematics 11c
  • 185247 09/25/06 18:00 Higinbotham Hall A Main Screen
  • 185238 09/25/06 17:05 A. Camsonne Beam calorimeter runs
  • 185229 09/25/06 16:27 A. Camsonne Gas Cerenkov PMT 6 HV increased
  • 185224 09/25/06 16:15 A. Camsonne portserver hatsv4 down
  • 185220 09/25/06 15:54 Ole Hansen Shift Summary
  • 185218 09/25/06 13:13 Elena Kh. Target screens
  • 185217 09/25/06 12:53 Ole Hansen HRSR magnets - recovering
  • 185216 09/25/06 12:02 jpchen target: HP stalled, reboot ioc, re-adjust, all fine now.
  • 185215 09/25/06 11:40 Ole Hansen Magnet crash
  • 185214 09/25/06 11:38 Ole Hansen Target problems
  • 185203 09/25/06 09:10 mpaolone incorrect momentum for runs 5179-5180
  • 185202 09/25/06 08:57 Ole Hansen HRSL angle off!
  • 185195 09/25/06 08:38 Elena Kh. LD2, 4 cm target
  • 185188 09/25/06 08:06 gron OWL Shift Summary
  • 185186 09/25/06 07:55 folts target chamber post
  • 185183 09/25/06 06:22 kelleher target motion: C
  • 185180 09/25/06 05:48 kelleher target screens
  • 185178 09/25/06 05:23 gron Moved to Kin 11
  • 185176 09/25/06 05:12 kelleher target motion: Ta
  • 185163 09/25/06 04:11 kelleher target motion: H2
  • 185156 09/25/06 03:00 kelleher target motion: 4cm Dummy
  • 185137 09/25/06 00:57 kelleher target motion: 4cm D2
  • Sunday

  • 185129 09/24/06 23:50 hyde 9/24/06 Swing Summary
  • 185124 09/24/06 23:27 A. Camsonne Target screens
  • 185121 09/24/06 23:15 A. Camsonne Target moved to C
  • 185118 09/24/06 22:35 hyde Kin 10A
  • 185111 09/24/06 21:59 A. Camsonne Target moved to Tantalum
  • 185098 09/24/06 20:31 A. Camsonne Target moved to C target
  • 185093 09/24/06 20:15 A. Camsonne Target moved to 4 cm dummy
  • 185090 09/24/06 20:04 hyde BCM Strip chart
  • 185077 09/24/06 19:05 hyde Vertical Steering
  • 185076 09/24/06 19:02 hyde Screen Shot of General Tools
  • 185075 09/24/06 19:01 A. Camsonne Target screens
  • 185064 09/24/06 18:22 A. Camsonne Left Arm to 25.5 degrees for Kin #9
  • 185038 09/24/06 16:04 A. Camsonne Target moved to 4 cm LD2
  • 185033 09/24/06 15:50 jglister Target to 4cm Dummy
  • 185032 09/24/06 15:49 Ransome shift summary
  • 185019 09/24/06 14:45 jglister Target to carbon
  • 185006 09/24/06 14:03 jglister Target to tantalum
  • 184993 09/24/06 13:07 jglister Target to carbon
  • 184988 09/24/06 12:54 jglister Target to 4cm Dummy
  • 184975 09/24/06 11:54 jglister Target to 4cm LD2
  • 184974 09/24/06 11:52 kuchina BPM strip charts screen
  • 184973 09/24/06 11:51 kuchina Hall A General Tools snapshot
  • 184970 09/24/06 11:43 jglister Target screens
  • 184961 09/24/06 10:59 Ransome L-dipole problem
  • 184954 09/24/06 10:47 ransome screen shot controls
  • 184951 09/24/06 08:59 kuchina Wrong LHRS momentum entry for runs 5124-5126, 25124-25126
  • 184946 09/24/06 08:04 gron Owl Shift Summary
  • 184938 09/24/06 06:04 kelleher target screens capture
  • 184937 09/24/06 06:00 kelleher target motion: LH2
  • 184932 09/24/06 05:07 gron Switched to Kin 6c
  • 184927 09/24/06 04:17 kelleher target motion: 4cm Al dummy
  • 184916 09/24/06 00:59 kelleher target: target screen oddities
  • 184915 09/24/06 00:53 byung dp with cut y for C
  • 184914 09/24/06 00:33 kelleher target motion: LD2
  • 184911 09/24/06 00:06 gron Switched to kinematics 6b
  • 184908 09/24/06 00:04 hyde Shift Summary Swing 09-23
  • Saturday

  • 184906 09/23/06 23:20 A. Camsonne Target moved to C
  • 184901 09/23/06 22:57 A. Camsonne Target moved to 4cm dummy
  • 184898 09/23/06 22:27 hyde HRS-L problem?
  • 184893 09/23/06 21:34 A. Camsonne Target moved to C
  • 184890 09/23/06 21:08 hyde Dipole He Supply
  • 184889 09/23/06 21:07 hyde Hall A Tools
  • 184888 09/23/06 21:03 hyde Hall A Tools
  • 184887 09/23/06 20:59 byung wrong momentum.
  • 184885 09/23/06 20:51 A. Camsonne Target moved to Ta
  • 184875 09/23/06 19:55 A. Camsonne Run 5110 Deuterium quasi-elastic
  • 184864 09/23/06 19:14 A. Camsonne Target moved to LH2 4cm
  • 184857 09/23/06 18:34 A. Camsonne Target moving to LD2
  • 184842 09/23/06 17:15 A. Camsonne Target screens flows
  • 184841 09/23/06 17:14 A. Camsonne Target screens
  • 184840 09/23/06 17:01 A. Camsonne Pedestals for Gas Cerenkov updated
  • 184835 09/23/06 16:31 A. Camsonne Target screens 4 cm LD2
  • 184833 09/23/06 15:59 Ransome Shift summary
  • 184828 09/23/06 15:45 kuchina BPM strip chart screen snapshot
  • 184827 09/23/06 15:17 kuchina Ta target run#5102(#25102)
  • 184826 09/23/06 15:14 Jiang and Le Looked at T3 and T4 events in run 5102
  • 184819 09/23/06 14:28 jglister Target to 4cm LD2
  • 184814 09/23/06 13:40 jglister Target to Tantalum
  • 184811 09/23/06 13:30 kuchina HallA General Tools Snapshot
  • 184810 09/23/06 13:11 jglister Target to 4cm Dummy
  • 184797 09/23/06 12:20 jglister Target screens
  • 184796 09/23/06 11:52 ransome high dead time
  • 184785 09/23/06 10:44 jglister Target to Carbon
  • 184768 09/23/06 09:25 jglister Target to 4cm LH2
  • 184759 09/23/06 08:38 jglister Target to 4cm Dummy
  • 184756 09/23/06 08:04 R. Michaels didn't cycle quads (should ?)
  • 184754 09/23/06 08:00 R. Michaels owl summary
  • 184745 09/23/06 06:44 kelleher target screen
  • 184742 09/23/06 06:35 kelleher target motion: deuterium
  • 184737 09/23/06 06:13 kelleher target motion: 4cm dummy
  • 184736 09/23/06 06:12 R. Michaels change P0L to 0.63278
  • 184735 09/23/06 06:08 R. Michaels spot++ run *85, next: 4cm dummy
  • 184726 09/23/06 05:49 R. Michaels spot, run 5083
  • 184721 09/23/06 05:20 kelleher target motion: C
  • 184718 09/23/06 05:08 Elena Kh. current in #5082, #25082
  • 184713 09/23/06 04:50 R. Michaels Qasy and scaler check
  • 184710 09/23/06 03:44 kelleher target motion: Tantalum
  • 184709 09/23/06 03:39 R. Michaels move to kine 4
  • 184700 09/23/06 02:15 kelleher no beam: RF separation
  • 184694 09/23/06 01:48 kelleher target motion: 4cm H2
  • 184693 09/23/06 01:41 R. Michaels kine 3c
  • 184680 09/23/06 00:17 R. Michaels tools screen / status
  • 184679 09/23/06 00:17 kelleher target motion: 4cm Dummy
  • Friday

  • 184676 09/22/06 23:58 gilman 30.5 deg for hrs left in kin 3
  • 184675 09/22/06 23:49 Ransome Shift summary
  • 184652 09/22/06 21:09 gron Moving target to LD2
  • 184593 09/22/06 16:05 LeRose Shift Summary
  • 184583 09/22/06 15:35 LeRose Target 4cm Deuterium
  • 184578 09/22/06 15:22 kuchina wrong left arm angle for runs #5048,#25048,#5049,#25049
  • 184563 09/22/06 10:48 LeRose Beam Permit
  • 184562 09/22/06 10:39 LeRose Power Permit
  • 184561 09/22/06 09:34 LeRose Hall A Tools
  • 184560 09/22/06 08:48 LeRose Controlled Access
  • 184558 09/22/06 07:57 gron Owl Shift Summary
  • 184557 09/22/06 07:21 gron No beam till 10:30
  • 184556 09/22/06 04:56 gron MCC
  • 184555 09/22/06 04:55 Elena Kh. wrong l.angle in #5047,#25047,#5046,#25046
  • 184554 09/22/06 04:44 Adaq target screen captured
  • 184553 09/22/06 04:43 gron MCC
  • 184552 09/22/06 04:08 gro MCC
  • 184551 09/22/06 04:07 kelleher target motion: carbon target
  • 184542 09/22/06 03:44 kelleher target motion: tantalum
  • 184541 09/22/06 03:43 gron kinematics
  • 184536 09/22/06 03:15 gron MCC
  • 184535 09/22/06 02:53 gron MCC
  • 184534 09/22/06 02:45 kelleher target motion: optics
  • 184529 09/22/06 02:19 kelleher target motion: carbon target
  • 184524 09/22/06 02:04 gron collimator
  • 184521 09/22/06 01:43 kelleher target heater setting
  • 184518 09/22/06 01:39 kelleher target motion: Al dummy
  • 184513 09/22/06 01:11 gron MCC
  • 184512 09/22/06 00:39 Adaq MCC
  • 184502 09/22/06 00:06 kelleher target motion: LH2
  • Thursday

  • 184499 09/21/06 23:59 slifer shift summary
  • 184492 09/21/06 23:39 slifer target fan speed increased
  • 184487 09/21/06 23:15 slifer target to 4cm LD2
  • 184482 09/21/06 22:25 slifer Target motion (and error)
  • 184459 09/21/06 20:43 R. Michaels Qasy feedback turned on
  • 184448 09/21/06 20:09 A. Camsonne Final tune for S1
  • 184432 09/21/06 19:22 A. Camsonne Pedestals for S1 updated in database
  • 184410 09/21/06 18:29 A. Camsonne S1 amplitude and HV
  • 184399 09/21/06 17:42 A. Camsonne Right arm DAQ stuck
  • 184374 09/21/06 16:25 slifer beam back to hall
  • 184370 09/21/06 15:52 Higinbotham Shift Summary
  • 184357 09/21/06 13:59 Higinbotham DAQ Fixed, but now no beam
  • 184348 09/21/06 13:34 A. Camsonne S1 not plugged correctly at EDTM module
  • 184347 09/21/06 13:27 atrig Trigger-Download
  • 184346 09/21/06 12:07 Higinbotham Controlled Access
  • 184343 09/21/06 11:45 atrig Trigger-Download
  • 184341 09/21/06 11:42 Higinbotham T1 vs. T3 Rate Asymmetry
  • 184339 09/21/06 11:17 Higinbotham Harp Check of Beam
  • 184338 09/21/06 11:09 gilman target problem
  • 184336 09/21/06 10:21 Fansler Left Q2 Ramp Problem
  • 184335 09/21/06 08:49 Higinbotham Tech: Left Q2 Trips Again
  • 184334 09/21/06 08:09 Higinbotham Target Status
  • 184331 09/21/06 07:19 X. Jiang left arm Q2 fault
  • 184316 09/21/06 06:39 gilman beam offered
  • Wednesday

  • 184313 09/20/06 22:50 Ramesh "VESDA" beeping
  • 184312 09/20/06 17:44 Ramesh Target Charts
  • 184311 09/20/06 17:34 Ramesh Target loop-1 repaired
  • 184310 09/20/06 16:24 folts tech/target chamber
  • 184309 09/20/06 15:34 M Jones Replaced -5V fuse in FPP Ch1 U plane M10
  • 184306 09/20/06 11:54 A. Camsonne FPP hot channel
  • Tuesday

  • 184304 09/19/06 15:14 gilman radiator
  • Sunday

  • 184292 09/17/06 15:52 R. Michaels T3 is clean, definitions of T4,T6
  • 184287 09/17/06 14:21 atrig Trigger-Download
  • 184267 09/17/06 13:44 atrig Trigger-Download
  • 184266 09/17/06 13:34 atrig Trigger-Download
  • 184265 09/17/06 13:33 atrig Trigger-Download
  • 184264 09/17/06 13:24 atrig Trigger-Download
  • 184263 09/17/06 13:23 atrig Trigger-Download
  • 184262 09/17/06 13:22 atrig Trigger-Download
  • 184261 09/17/06 13:11 atrig Trigger-Download
  • 184260 09/17/06 13:11 atrig Trigger-Download
  • 184259 09/17/06 12:00 atrig Trigger-Download
  • 184258 09/17/06 11:47 atrig Trigger-Download
  • 184257 09/17/06 11:38 atrig Trigger-Download
  • 184256 09/17/06 11:17 atrig Trigger-Download
  • 184255 09/17/06 11:17 atrig Trigger-Download
  • 184254 09/17/06 11:17 atrig Trigger-Download
  • 184253 09/17/06 11:12 R. Michaels T4 mlu files corrected
  • 184250 09/17/06 10:37 R. Michaels note regarding T4
  • 184249 09/17/06 10:33 atrig Trigger-Download
  • 184248 09/17/06 10:31 atrig Trigger-Download
  • Friday

  • 184247 09/15/06 16:51 R. Michaels L-arm TC moved, ok now
  • 184245 09/15/06 14:59 R. Michaels Coinc. time
  • 184196 09/15/06 13:40 atrig Trigger-Download
  • 184195 09/15/06 13:33 atrig Trigger-Download
  • 184183 09/15/06 11:40 atrig Trigger-Download
  • Thursday

  • 184175 09/14/06 17:38 atrig Trigger-Download
  • 184174 09/14/06 17:26 atrig Trigger-Download
  • 184173 09/14/06 17:02 atrig Trigger-Download
  • 184165 09/14/06 13:42 gilman FPP
  • Wednesday

  • 184153 09/13/06 15:04 atrig Trigger-Download
  • 184152 09/13/06 11:51 R. Michaels strange T8 removed
  • 184151 09/13/06 11:19 atrig Trigger-Download
  • 184150 09/13/06 11:18 atrig Trigger-Download
  • 184148 09/13/06 10:49 atrig Trigger-Download
  • 184147 09/13/06 10:43 atrig Trigger-Download
  • Monday

  • 184145 09/11/06 15:02 atrig Trigger-Download
  • 184135 09/11/06 08:39 folts ESR follow up
  • Friday

    
    

    Thursday

  • 184121 09/07/06 15:43 atrig Trigger-Download
  • 184118 09/07/06 13:01 folts tech/ESR
  • Wednesday

    
    

    Tuesday